# MuSC: Improving Complex Instruction Following with Multi-granularity Self-Contrastive Training

### v2

@@ 1 Introduction
Large Language Models (LLMs) have made remarkable advancements and are being wildly applied across various domains Zhao et al. (2024); Minaee et al. (2024); Hadi et al. (2024); Yin et al. (2024); Wang et al. (2024a). The instruction-following ability is fundamental and important, as it enables LLMs to generate appropriate responses to given instructions and solve corresponding tasks OpenAI et al. (2024). While recent LLMs perform comparatively well on simple instructions, their response quality to complex instructions with elaborate constraints often falls under expectation, with some of the constraints omitted He et al. (2024); Jiang et al. (2024b), which hinders their application in more real-world complex scenarios.
@@
Recently, the research community has paid attention to self-alignment, to break the data bottleneck without relying on a stronger model Wang et al. (2024b); Zhang et al. (2024a). Self-Reward Yuan et al. (2024) proposes to utilize the model itself to both generate responses and evaluate the results, which can be incorporated into DPO training. ISHEEP Liang et al. (2024) proposes an automatic loop to self-assess and self-filter instruction data. Despite their effectiveness, these methods are targeted at general instruction-following ability. The self-alignment of complex instruction-following ability remains unexplored.
@@ 2 Related Work
Complex Instruction-Following. As one of the cores of LLM intelligence, how to improve the model’s instruction-following capability is important. The earliest works, such as Alpaca Taori et al. (2023), Vicuna Chiang et al. (2023), and Camel Wang et al. (2023), used instruction data generated by proprietary models to supervise fine-tuning of open-source models, significantly enhancing their instruction-following capabilities. However, these methods mainly focus on general instruction following, while complex instruction following still remains challenging. To cope with this challenge, a lot of methods Yin et al. (2023); Lou et al. (2023); He et al. (2024); Sun et al. (2024b); Chen et al. (2024b); Dong et al. (2024) have been proposed to construct complex instruction data. The earliest work is Evol-Instruct Xu et al. (2024), which proposed to utilize GPT-4 to expand the instructions from both depth and width, thereby generating complex instructions and corresponding constraints. Conifer Sun et al. (2024a) proposed a progressive learning strategy designed to help smaller models incrementally enhance their abilities.
@@
Self-Alignment. Self-alignment refers to aligning the model to human preference without relying on a more advanced model or external supervision. As an early study, Self-Rewarding Yuan et al. (2024) proposed the model itself to both generate responses and evaluate the results. Following this work, many works Liu et al. (2024); Chen et al. (2024b, a); Pang et al. (2024); Meng et al. (2024) are conducted to obtain supervision data by the model itself. Meta-Rewarding Wu et al. (2024) advanced the concept by improving the model’s instruction and evaluation capabilities simultaneously. Liu et al. (2024) employed diverse prompts to guide the model to generate various responses. Despite the progress these methods have made, they all target general instruction following. For complex instructions with multiple constraints, the response will be lengthy and multi-facet, resulting in challenges for the self-evaluation process.
@@ 5.2 The Influence of Confidence Metrics
As shown in Table 3, entropy-based token weight achieves the best result among all metrics, verifying its effectiveness. Both the perplexity and PMI-based score underperforms, as they only consider the probability of the selected token instead of the whole distribution, leading to biased evaluation. KLDiv-based score also underperforms, this is because KLDiv is essentially a distance measurement instead of a confidence measurement, which is not adapted to our scenario.
@@ 5.4 Is Fundamental Capability Harmed?
Table 5: Experiment results of our proposed methods on fundamental capability benchmarks.
@@ Appendix D Detailed Experiment Results
In this section, we presented detailed experiment results which are omitted in the main body of this paper due to space limitation. The detailed experiment results of different methods on ComplexBench, FollowBench and AlpacaEval2 are presented in Table 7, 9 and 8. The detailed results for the ablative studies of confidence metrics is presented in Table 10. The detailed results for the ablative studies of confidence metrics is presented in Table 6. We also present a case study in Table 11, which visualize the token-level weight derived from calibrated confidence score.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have made remarkable advancements and are being wildly applied across various domains Zhao et al. (2024); Wu et al. (2024a); He et al. (2024c); Liu et al. (2025); He et al. (2024b); Zhang et al. (2025). The instruction-following ability is fundamental and important, as it enables LLMs to generate appropriate responses to given instructions and solve corresponding tasks OpenAI et al. (2024). While recent LLMs perform comparatively well on simple instructions, their response quality to complex instructions with elaborate constraints often falls under expectation, with some of the constraints omitted He et al. (2024a); Jiang et al. (2024b), which hinders their application in more real-world complex scenarios.
@@
Recently, the research community has paid attention to self-alignment, to break the data bottleneck without relying on a stronger model Wang et al. (2024); Zhang et al. (2024a). Self-Reward Yuan et al. (2024) proposes to utilize the model itself to both generate responses and evaluate the results, which can be incorporated into DPO training. ISHEEP Liang et al. (2024) proposes an automatic loop to self-assess and self-filter instruction data. Despite their effectiveness, these methods are targeted at general instruction-following ability. The self-alignment of complex instruction-following ability remains unexplored.
@@ 2 Related Work
Complex Instruction-Following. As one of the cores of LLM intelligence, how to improve the model’s instruction-following capability is important. The earliest works, such as Alpaca Taori et al. (2023), Vicuna Chiang et al. (2023), and Camel Wang et al. (2023), used instruction data generated by proprietary models to supervise fine-tuning of open-source models, significantly enhancing their instruction-following capabilities. However, these methods mainly focus on general instruction following, while complex instruction following still remains challenging. To cope with this challenge, a lot of methods Yin et al. (2023); Lou et al. (2023); He et al. (2024a); Sun et al. (2024b); Chen et al. (2024b); Dong et al. (2024) have been proposed to construct complex instruction data. The earliest work is Evol-Instruct Xu et al. (2024), which proposed to utilize GPT-4 to expand the instructions from both depth and width, thereby generating complex instructions and corresponding constraints. Conifer Sun et al. (2024a) proposed a progressive learning strategy designed to help smaller models incrementally enhance their abilities.
@@
Self-Alignment. Self-alignment refers to aligning the model to human preference without relying on a more advanced model or external supervision. As an early study, Self-Rewarding Yuan et al. (2024) proposed the model itself to both generate responses and evaluate the results. Following this work, many works Liu et al. (2024); Chen et al. (2024b, a); Pang et al. (2024); Meng et al. (2024) are conducted to obtain supervision data by the model itself. Meta-Rewarding Wu et al. (2024b) advanced the concept by improving the model’s instruction and evaluation capabilities simultaneously. Liu et al. (2024) employed diverse prompts to guide the model to generate various responses. Despite the progress these methods have made, they all target general instruction following. For complex instructions with multiple constraints, the response will be lengthy and multi-facet, resulting in challenges for the self-evaluation process.
@@ 5.2 The Influence of Confidence Metrics
As shown in Table 3, entropy-based token weight achieves the best result among all metrics, verifying its effectiveness. Both the perplexity and PMI-based score underperforms, as they only consider the probabilities of the selected tokens instead of the whole distribution, leading to biased evaluation [^6] As the model would always select (one of) the tokens with the highest probability at each step, relying only on the the selected tokens for evaluation will result in overconfidence.. KLDiv-based score also underperforms, this is because KLDiv is essentially a distance measurement instead of a confidence measurement, which is not adapted to our scenario.
@@ 5.3 Can MuSC Scale to Other xPO Method?
While IPO showed the best performance among different xPO methods, these performance differences were not statistically significant. Our method is not limited to DPO; instead, it’s a general framework applicable to most xPO-style methods. Given the popularity of DPO, we primarily report our experiments based on DPO in this work.
@@ 5.4 Is Fundamental Capability Harmed?
Table 5: Experiment results of our methods on fundamental capability benchmarks. Bolded results denote improvements, while grayed results denote degradation.
@@ 5.5 The Variation of Statistical Indicators
5.6 Can MuSC Scale among Model Sizes
Model CFBench AlpacaEval2
CSR ISR PSR LC Len
Qwen2.5-1.5B-Inst 0.59 0.21 0.28 10.10 1507
+MuSC 0.69 0.32 0.41 11.44 1346
Qwen2.5-3B-Inst 0.72 0.34 0.45 27.44 2188
+MuSC 0.76 0.41 0.52 30.64 1873
Qwen2.5-14B-Inst 0.82 0.49 0.60 46.05 1792
+MuSC 0.85 0.56 0.67 47.61 1770
Qwen2.5-32B-Inst 0.87 0.60 0.70 48.63 1785
+MuSC 0.89 0.65 0.75 49.34 1772
Table 6: Experiment results of MuSC on Qwen2.5-Instruct of different model sizes.
To verify the scalability of our proposed method on models with larger sizes, we conducted extensive experiments on Qwen-2.5-Instruct models ranging from 1.5B to 32B parameters. As shown in Table 6, the results demonstrate consistent improvements across both smaller and larger models, confirming that our method is effective regardless of model size. This robust scalability is attributed to our approach being inherently self-improving and not requiring external models, which holds significant promise for advancing LLM capabilities.
@@ 6 Conclusion
Ethical Considerations
Since the aligned data is generated based on the Instruct version model, the chances of generating toxic content are low. To further validate this, we randomly selected 200 samples in total—100 samples for Llama-3-8B-Instruct and 100 samples for Qwen-2-7B-Instruct—and employed three graduate students to annotate them. The results showed no instances of toxic instructions or responses.
However, there is potential risk of knowledge forgetting related to safety alignment. To address this, we recommend incorporating additional safe alignment samples during MuSC training to reinforce the model’s safety-related knowledge. We will include these recommendations in the ethical considerations section of the revised manuscript.
Acknowledgements
This work is supported by National Natural Science Foundation of China (62276077, 62376075, 62376076), Jiangsu Science and Technology Major Project (BG2024031) and Nanjing University AI & AI for Science Funding (2024300540).
@@ Appendix D Detailed Experiment Results
In this section, we presented detailed experiment results which are omitted in the main body of this paper due to space limitation. The detailed experiment results of different methods on ComplexBench, FollowBench and AlpacaEval2 are presented in Table 8, 10 and 9. The detailed results for the ablative studies of confidence metrics is presented in Table 11. The detailed results for the ablative studies of confidence metrics is presented in Table 7. We also present a case study in Table 12, which visualize the token-level weight derived from calibrated confidence score.